# Non-local, diamagnetic electromagnetic effects in long MITLs

### v1

@@ III Non-local, diamagnetic electromagnetic effects of SCL currents
We observe from Eqs. (6), (7) that in a long MITL, a wave emitted by a local SCL current per the “once-on-always-on” rule propagates throughout the full length of the MITL with undiminished amplitude, i.e., is nonlocal. The diamagnetic nature of the waves can be seen from Fig. 1. The figure also shows the correspondence between the SCL current within the volume of the MITL, $\mathbf{j}_{\rm SCL}$, and the changes in the currents on the electrodes, namely, $I_{0}=j_{\rm SCL}/2$. We discuss these two effects below.
@@
Figure 1: Directions of fields and currents within the MITL and on the conductor. Currents with subscript ’0’ denote current on the conductor. The field directions are shown only for waves propagating in the positive $z$-direction; waves propagating to the left have reversed magnetic field direction. Both self-waves are thus diamagnetic.
@@
The non-local effect of SCL currents is demonstrated in Fig. 2, where the temporal evolution of a wave from a point current source at $z=50\,$cm with $I_{0}=100\,$kA is shown [cf. Eq. (1)]. The current source indicated in the figure is in the negative $x$-direction, hence, the fields have opposite signs to those in Eqs. (6), (7). The MITL has a length of $L=1\,$m and an AK gap $d=5\,$mm. At $t=0.5\,$ns, two waves are seen propagating away from the source (directions are indicated by arrows) and before reflection from the short at $z=1\,$m. At $t=2.5\,$ns, the right-propagating wave has reflected and is now propagating to the left, while the left-propagating wave has exited the MITL. It is seen that since the amplitudes of the waves due to this source are non-diminishing in time and in space (up to the wave front), after reflection the electric field of the right-propagating wave cancels out exactly the electric field of the reflected (now left-propagating) wave so that the total electric field behind the reflected wave is exactly zero. The magnetic fields of the incoming and outgoing waves add and at $t=5.5\,$ns a steady state has been established with only a constant magnetic field present in the MITL. The jump in the magnetic field can be calculated from Eq. (8) and equals $\left[B_{y}\right]_{z=50\,\mbox{cm}}\simeq 0.126\,$T. In a time-dependent system (time-dependent current source), the emitted and reflected waves would not cancel exactly due to the temporal delay between the time of emission and the time the reflected wave has traveled back to the source, at which time the emitted wave from the source will have changed its amplitude. Note that this relatively low current (peak currents of interest in the $Z$ machine are 20202020303030\,30MA, while smaller pulsed power machines have currents of order mega-ampere) already generates an electric field with amplitude $\simeq 19\,$MV/m, which is very close to the threshold for SCL emission discussed in the Introduction. Conversely, currents of this magnitude may be expected during the beginning of SCL emission.
@@
The large diamagnetic response of the SCL currents is demonstrated in Fig. 3 at three times, 6.56.56.5\,6.5ns, 999\,9ns, and 242424\,24ns. For this setup, we use a sine squared drive current pulse with $I_{\rm peak}=20\,$MA and $\tau_{\rm peak}=100\,$ns. SCL emission starts at about 555\,5ns. In about 1.51.51.5\,1.5ns, the electric field measured near the anode has accumulated an amplitude comparable to the threshold for SCL emission, 242424\,24MV/m (top panel). The middle panel of the figure shows the diamagnetic wave after reflection from the conducting short at $z=1\,$m. The bottom panel shows the electromagnetic fields after magnetic insulation, which approach the vacuum field values (black dash-dotted lines). The similarity between the fields in Figs. 3 and 2 is obvious with the difference that in the time-dependent case, the large diamagnetic EM field is generated by an SCL current that has accumulated in time (and is spread out in space).
@@
Another typical consequence of the diamagnetic effect is shown in Fig. 4, where current densities at the anode along the extent of the MITL are shown. Non-zero SCL current densities are also an indication of where the SCL emission threshold has been exceeded. The top panel shows that in about 1.51.51.5\,1.5ns of (uninsulated) SCL emission, about 303030\,30cm of cathode (surface) extent has started emitting according to the PIC model and about 353535\,35cm in the 1D model. At time 999\,9ns, shown in the bottom panel, since the diamagnetic EM fields subtract from the vacuum fields, the total resulting electric field in the remaining part of the MITL extent ($\gtrsim 30\,$cm) has fallen below the emission threshold, halting SCL emission almost completely for the next 333\,3ns. The vacuum fields at 999\,9ns are shown in the middle panel of Fig. (3) and suggest that without this diamagnetic effect, SCL emission would have initiated in 0similar-toabsent0$\sim$ 0∼ 0606060\,60cm of the cathode surface.
This dynamics of fields and currents tends to repeat itself and thus the emitting fraction of MITL surface grows in “jumps” instead of gradually, as one might expect from a smooth temporal drive. In other words, for the time interval before the full cathode surface starts emitting (this is roughly the time for the first diamagnetic wave to reach the load), there is a sequence of a propagating SCL current pulse followed by a pause. During the propagating part of the cycle, the diamagnetic effect accumulates until the external wave’s amplitude falls bellow the SCL emitting threshold. During the pause part of the cycle, the external fields grow sufficiently to eventually overwhelm the diamagnetic effect.
@@
We see that the non-local diamagnetic response acts to lower the magnitude of SCL currents and reduce current losses in two ways. First, by limiting the extent of emitting surface. This effect varies in magnitude in various setups, however, it can be used as a telltale sign of the non-local, diamagnetic response of SCL currents. And second, by lowering the SCL current within the emitting surface (recall $j_{\rm SCL}\sim E^{3/2}$). This effect is bi-directional since local SCL currents emit diamagnetic EM waves in both propagating directions. The result of this is observed in Fig. 4 (bottom panel) where the current density near the launch side ($z=0$) has been greatly reduced, forming a sharp peak near $z\simeq 30\,$cm. It is the combination of these two effects that results in dramatically lower overall current losses in long MITLs compared to SCL emission in vacuum fields alone (see also the discussion of Fig. 5 below).
Some interesting observations can relate our results to previous work. First, the peak of the uninsulated current density propagates about 232323\,23cm in 1.51.51.5\,1.5ns, which is approximately at the speed of $0.65c$. The current pulse observed (on the conductor) in the MITL is the difference between the vacuum current and the loss current, and it would therefore appear that the main current pulse propagates at that speed as well. Second, since the the SCL (loss) current is seen to have a sharp peak, the main current pulse would exhibit “front sharpening.” Both features have been previously reported by experimental and numerical work baranchikov_transfer_1977 ; di_capua_propagation_1979 ; bergeron_equivalent_1977 . On the other hand, the “jump-like” propagation discussed above has not been previously reported.
@@ IV Scaling studies of MITL losses based on the 1D electromagnetic model
We have tested relatively extensively the validity of the 1D model against PIC simulations, varying parameters such as MITL dimensions (length, AK gap), peak current, pulse length, as well as threshold for SCL emission within relevant ranges. In very short MITLs, for example, lowering $E_{\rm th}$ to artificially low values is necessary in order to observe uninsulated SCL emission.
@@
where $j_{\rm SCL}(z,t)$ is the uninsulated current and note that it is a function of time alone. A comparison of predicted current loss by the 1D model as well as EMPIRE and CHICAGO 2D PIC is shown in Fig. 5 (left-hand scale). We note the excellent agreement between all three codes. The input current pulse (without SCL emission) is indicated with a dash-dotted black line. Also plotted in Fig. 5 is the loss current (by the 1D model) based on only the Child-Langmuir law in vacuum fields (right-hand scale), i.e., without accounting for diamagnetic EM fields emitted by SCL currents. We see that such a prediction, even for a 111\,1m long MITL, is grossly inaccurate, with a maximum current loss of over 888\,8MA compared to the expected 250250250\,250kA. (The fact that in this case the loss current exceeds the input current should be no cause for alarm. It is an artifact of the simulation setup used throughout this work, where we use a feed of infinite source of energy. Using such a feed removes unnecessary complications due to variations in the input drive current and allows for a clean demonstration of the physics under consideration.)
Figure 5: Current loss vs. time: Comparison between the 1D model and 2D PIC simulations with EMPIRE and CHICAGO (left-hand scale). The dashed blue curve (right-hand scale) shows the current loss without accounting for non-local diamagnetic effects.
@@
and the time integral is until the time $t_{\rm ins}$ of complete magnetic insulation along the MITL; then $Q_{\rm loss}$ is the total charge collected by the anode. We normalize this to the total charge delivered by the drive current, which for a sine squared pulse equals $Q_{\rm tot}=I_{\rm peak}\tau_{\rm peak}$.
@@
Figure 6: Scaling studies of loss charge. Panel (a): Scaling with AK gap, $d$, for fixed $I_{\rm peak}=20\,$MA and $\tau_{\rm peak}=100\,$ns. Panel (b): Scaling with peak current, $I_{\rm peak}$, for a fixed AK gap, $d=5\,$mm and $\tau_{\rm peak}=100\,$ns. Panel (c): Scaling with pulse length, $\tau_{\rm peak}$, for a fixed $I_{\rm peak}=20\,$MA and AK gap, $d=5\,$mm. Panel (d): Scaling with MITL length, $L$, for a fixed $I_{\rm peak}=20\,$MA and $\tau_{\rm peak}=100\,$ns.
@@
In Fig. 6 (a) we show the scaling of MITL losses with the AK gap distance, $d$, for three different lengths. The pulse has $I_{\rm peak}=20\,$MA and $\tau_{\rm peak}=100\,$ns. We observe that for small AK gap sizes, $Q_{\rm loss}/Q_{\rm tot}$ scales as a power law $\sim 1/x$ while for larger gap sizes that scaling does not hold. In practical applications, small gap sizes are usually preferred and then the $\sim 1/x$ scaling can be used.
@@
The next Fig. 6 (b) shows how MITL losses scale with peak current, $I_{\rm peak}$, for a given fixed MITL length, pulse length $\tau_{\rm peak}=100\,$ns, and $d=5\,$mm. The fractional losses exhibit a maximum at a certain value of peak current, vanish for small currents, and decrease for large peak currents. We can understand this behavior in the following way. On the one hand, smaller peak currents have smaller losses due to the shorter time of uninsulated SCL emission: regardless of the drive current amplitude, SCL emission turns on only after exceeding some threshold value, say $E_{\rm th}=24\,$MV/m in our case. If the peak current is too low, that threshold value cannot be exceeded since for a fixed $\tau_{\rm peak}$, $E\sim I_{\rm peak}$. On the other hand, large peak currents induce larger diamagnetic effects, which lower the losses. Obviously, peak currents around the maximum of the curve should be avoided in MITL designs. A common scaling of MITL losses for large peak currents emerges in the form of $\sim 1/x^{0.8}$, which we leave as an empirical observation.
@@
In Fig. 6 (c) we show the scaling of MITL losses with MITL length for fixed $I_{\rm peak}=20\,$MA and $\tau_{\rm peak}=100\,$ns, for three different AK gap sizes. At shorter lengths, the current losses drop due to the decreased time of uninsulated SCL emission. In the limit of (infinitely) short MITLs, we converge to the losses predicted by the Child-Langmuir law in vacuum fields (not shown). Similarly, we can understand the increased current losses in longer MITLs as increased time of uninsulated SCL emission. Recall that magnetic insulation cannot happen in pure TEM fields, i.e., for which $E=cB$. Electrons perform a figure-$8$ motion in such fields, not a cyclotron motion, and our estimates have shown that the size of such an orbit for typical fields of interest is much larger than the AK gap; this means such electrons are uninsulated. Magnetic insulation can occur after reflection from the load because the relation $E=cB$ no longer holds due to the interference (superposition) of the incoming and outgoing waves. In longer MITLs, reflection takes longer to happen, hence, the increased losses. For the range of lengths shown, we again uncover an empirical scaling law for large $L$, which, however, differs for the different gap sizes.
@@
It is worth commenting on the losses in infinitely long MITLs. Although magnetic insulation never occurs in such MITLs, current losses are still limited by two effects: (i) the Child-Langmuir law limits the amount of current grossing the AK gap; and (ii) the non-local diamagnetic effects further lower the amount of loss current; in fact, these latter effects are maximally manifested in this case.
@@
The last Fig. 6 (d) shows scaling of MITL losses with pulse duration. This scaling is the strongest of the four examples shown in the figure, showing exponential or super-exponential dependence (notice the semi-log scale of the plot). The losses for the 111\,1m-long MITL fall to zero past about $\tau_{\rm peak}\gtrsim 300\,$ns. These losses can also be understood in terms of ununsulation time, which is longer for longer MITLs. Another intuitive understanding can be given by the relation $E\sim dI/dt$, showing lower fields for longer pulse lengths. (This relation strictly applies only in the infinitely short MITL or equivalently, the limit of $c\rightarrow\infty$, see also Ref. hess_electron_2021 ). No common empirical scaling law emerges in this case.

### v2

@@ III Non-local, diamagnetic electromagnetic effects of SCL currents
We observe from Eqs. (6), (7) that in a long MITL, a wave emitted by a local SCL current per the “once-on-always-on” rule propagates throughout the full length of the MITL with undiminished amplitude, i.e., is non-local. The diamagnetic nature of the waves can be seen from Fig. 1. The fields labeled “self” are emitted by SCL currents, while those labeled “ext” are the fields launched externally into the MITL (by a generator). The figure also shows the corresponding currents on the electrodes (conductors). In the rest of the paper we omit the “self” and “ext” subscripts when we see no confusion; we also use the terms “external” and “vacuum” fields interchangeably. We discuss these two effects next.
@@
Figure 1: Directions of fields and currents within the MITL and on the conductor. Currents with subscript ’0’ denote current on the conductor. The field directions are shown only for waves propagating in the positive $z$-direction; waves propagating to the left have a reversed magnetic field direction. The fields labeled “self” are emitted from the point source current $\mathbf{j}_{\rm SCL}$ [cf. Eq. (1)] whereas the fields labeled “ext” are launched by an external (generator) source [cf. Eqs. (9)–(11)]. The self-waves in both directions are thus diamagnetic.
@@
The non-local effect of SCL currents is demonstrated in Fig. 2, where the temporal evolution of a wave from a point current source at $z_{0}=50\,$cm with $I_{0}=100\,$kA/m is shown [cf. Eq. (1)]. The current source indicated in the figure is in the negative $x$-direction, hence, the fields have opposite signs to those in Eqs. (6), (7). The MITL has a length of $L=1\,$m and an AK gap $d=5\,$mm. At $t=0.5\,$ns, two waves are seen propagating away from the source (directions are indicated by arrows) and before reflection from the short at $z=1\,$m. At $t=2.5\,$ns, the right-propagating wave has reflected and is now propagating to the left, while the left-propagating wave has exited the MITL. It is seen that since the amplitudes of the waves due to this source are non-diminishing in time and in space (up to the wave front), the electric field of the emitted (right-propagating) wave cancels out exactly the electric field of the reflected (now left-propagating) wave so that the total electric field behind the reflected wave is exactly zero. The magnetic fields of the incoming and outgoing waves add, and at $t=5.5\,$ns, a steady state has been established with only a constant magnetic field present in the MITL. The magnitude of this magnetic field is zero for $z<0.5\,$m and jumps to $\left[B_{y}\right]_{z=50\,\mbox{cm}}\simeq 0.126\,$T for $z>0.5\,$m, according to Eq. (8). In a time-dependent system (time-dependent current source), the emitted and reflected waves would not cancel exactly due to the temporal delay between the time of emission and the time the reflected wave has traveled back to the source, during which the emitted wave from the source will have changed its amplitude. Note that this relatively low current (peak currents of interest in the $Z$ machine are 20202020303030\,30MA, while smaller pulsed power machines have currents of order mega-ampere) already generates an electric field with amplitude $\simeq 19\,$MV/m, comparable to the threshold for SCL emission. Conversely, currents of this magnitude may be expected during the beginning of SCL emission.
@@
The large diamagnetic response of the SCL currents is demonstrated in Fig. 3 at three times, 6.56.56.5\,6.5ns, 999\,9ns, and 242424\,24ns. For this setup, we use a sine squared drive current pulse with $I_{\rm peak}=20\,$MA/m and $\tau_{\rm peak}=100\,$ns. SCL emission starts at about 555\,5ns. At 6.56.56.5\,6.5ns, the electric field measured near the anode drops to near zero, indicating that the diamagnetic electric field amplitude is comparable to the threshold for SCL emission (top panel). (Unless specified otherwise, all simulations assume threshold for SCL emission 242424\,24MV/m.) The middle panel of the figure shows the diamagnetic wave after reflection from the conducting short at $z=1\,$m. The bottom panel shows the electromagnetic fields after magnetic insulation, which approach the vacuum fields (black lines). The bottom panel also shows that the electric field more closely approximates the vacuum field, while the magnetic field exhibits a small offset, differing by less than 1% near $z=1\,$m to a maximum of about 1.6% near $z=0\,$m. This offset can be explained by two observations. First, after magnetic insulation, the electrons form a layer near the cathode; that layer creates a magnetic field (also referred to as self-magnetic field in the literature), which our 1D model neglects. Second, numerically it is difficult to resolve the electromagnetic fields near the time of insulation (just before or shortly after) due to the fact that the statistics of computational particles reaching the anode becomes very poor, leading to increased numerical noise and subsequent accumulation of numerical errors that degrade the accuracy of the simulation.
We may note the similarity between the diamagnetic fields in Figs. 2 and 3, the difference being that in Fig. 3, the time-dependent external drive causes SCL currents to accumulate in time and to spread over a finite emitting cathode surface.
Another consequence of the diamagnetic effect is shown in Fig. 4, where current densities at the anode along the extent of the MITL are shown. Non-zero SCL current densities are also an indication of where the SCL emission threshold has been exceeded. The top panel shows that in about 1.51.51.5\,1.5ns of (uninsulated) SCL emission (we remind the reader that SCL emission initiates at about 555\,5ns), about 303030\,30cm of cathode (surface) extent has started emitting according to the PIC model and about 353535\,35cm in the 1D model. At time 999\,9ns, shown in the bottom panel, since the diamagnetic electric fields subtract from the vacuum fields, the total resulting electric field in the remaining part of the MITL extent ($\gtrsim 30\,$cm) has fallen below the emission threshold, halting SCL emission almost completely for the next 333\,3ns. The vacuum fields at 999\,9ns are shown in the middle panel of Fig. (3) and suggest that without this diamagnetic effect, SCL emission would have covered the extent of about 00606060\,60cm of cathode surface.
Such dynamics of fields and currents tends to repeat itself, causing the emitting fraction of MITL surface to grow in “jumps” instead of gradually, as one might expect from a smooth temporal drive. In other words, for the time interval before the full cathode surface starts emitting (this is roughly the time for the first diamagnetic wave to reach the load), there is a sequence of a propagating SCL current pulse followed by a pause. During the propagating part of the cycle, the diamagnetic effect accumulates until the total electric field amplitude falls bellow the SCL emitting threshold. During the pause part of the cycle, the external fields grow sufficiently so that the total field overwhelms the diamagnetic effect and the fraction of emitting surface extends further into the MITL.
@@
We see that the non-local diamagnetic response acts to lower current losses in two ways. First, by limiting the extent of emitting surface. This effect varies in magnitude in various setups, however, it can be used as a telltale sign of the non-local, diamagnetic response of SCL currents. And second, by lowering the electric field magnitude within the emitting surface, leading to lower SCL currents (recall $j_{\rm SCL}\sim E^{3/2}$). This latter effect is bi-directional since local SCL currents emit diamagnetic EM waves in both propagating directions. The result of this is observed in Fig. 4 (bottom panel) where the current density near the launch side ($z=0$) has been greatly reduced, forming a sharp peak near $z\simeq 30\,$cm. It is the combination of these two effects that causes dramatically lower current losses in long MITLs compared to losses by SCL emission in vacuum fields alone (see also the discussion of Fig. 5 below).
The bi-directional diamagnetic property also affects the magnetic fields within the MITL, as observed in Fig. (3). The top and middle panels of the figure show the magnetic fields in the MITL with and without SCL emission. We observe in both panels magnetic field enhancement (in absolute value) near $z=0$, while the middle panel also shows magnetic field reduction near $z=1\,$m. This is understood by referring to Fig. 1 and the top and middle panels of Fig. 2. Namely, we see that the magnetic field of an emitted to the left diamagnetic wave adds, while that of a diamagnetic wave emitted to the right (including after reflection) subtracts from the magnetic fields of the external wave.
Some interesting observations can relate our results to previous work. First, the peak of the uninsulated current density propagates about 232323\,23cm in 1.51.51.5\,1.5ns, which is approximately at the speed of $0.65c$. The current pulse observed (on the conductor) in the MITL is the difference between the vacuum current and the loss current, and it would therefore appear that the main current pulse propagates at that speed as well. Second, since the the SCL (loss) current is seen to have a sharp peak, the main current pulse would exhibit “front sharpening.” Both features have been previously reported by experimental and numerical work [4, 5, 40, 41]. On the other hand, the “jump-like” pulse propagation discussed above has not been previously reported.
@@ IV Scaling studies of MITL losses based on the 1D electromagnetic model
We have tested relatively extensively the validity of the 1D model against PIC simulations, varying parameters such as MITL dimensions (length, AK gap), peak current, pulse length, as well as threshold for SCL emission within relevant ranges. In tests with very short MITLs, for example, lowering $E_{\rm th}$ to artificially low values was necessary in order to observe uninsulated SCL emission.
@@
where $j_{\rm SCL}(z,t)$ is the uninsulated current and note that it is a function of time alone. A comparison of predicted current loss by the 1D model as well as EMPIRE and CHICAGO 2D PIC is shown in Fig. 5 (left-hand scale). We note the excellent agreement between all three codes. The input current pulse (without SCL emission) is indicated with a dash-dotted black line.
Also plotted in Fig. 5 is the loss current (by the 1D model) based on only space charge limiting effects, i.e., the Child-Langmuir law (right-hand scale). We see that such a prediction, even for a 111\,1m long MITL, is grossly inaccurate, with a maximum current loss of over 888\,8MA/m, compared to the expected 250250250\,250kA/m. That the loss current exceeds the input current should not be a cause for alarm. Such a setup is obviously not physical for the parameters at hand and in an electromagnetic simulation the loss current never exceeds the drive current. However, one may think of this as a quasistatic approximation, where the input drive changes so slowly that electromagnetic effects can be neglected; alternatively, one may think of this approximation as the limit $c\rightarrow\infty$. In this limit, the input drive acts as an ideal voltage source, which can provide arbitrarily large current to a circuit. Although not valid for the present case, the quasistatic approximation becomes valid in the limit of an infinitely short MITL. Indeed, we have verified that in this limit the quasistatic and electromagnetic simulations yield similar current losses (not shown).
Figure 5: Current loss vs. time: Comparison between the 1D model and 2D PIC simulations with EMPIRE and CHICAGO (left-hand scale). The dashed magenta curve (right-hand scale) shows the current loss without accounting for non-local diamagnetic effects.
@@
The integral’s upper limit is the time for complete magnetic insulation along the MITL, $t_{\rm ins}$, defined as $I_{\rm loss}(t)=0$ for $t>t_{\rm ins}$; then $Q_{\rm loss}$ is the total charge collected by the anode. We normalize this to the total charge delivered by the drive current, which for a sine squared pulse equals $Q_{\rm tot}=I_{\rm peak}\tau_{\rm peak}$.
@@
Figure 6: Scaling studies of loss charge. Panel (a): Scaling with AK gap, $d$, for fixed $I_{\rm peak}=20\,$MA/m and $\tau_{\rm peak}=100\,$ns. Panel (b): Scaling with peak current, $I_{\rm peak}$, for a fixed AK gap, $d=5\,$mm and $\tau_{\rm peak}=100\,$ns. Panel (c): Scaling with pulse length, $\tau_{\rm peak}$, for a fixed $I_{\rm peak}=20\,$MA/m and AK gap, $d=5\,$mm. Panel (d): Scaling with MITL length, $L$, for a fixed $I_{\rm peak}=20\,$MA/m and $\tau_{\rm peak}=100\,$ns.
@@
In Fig. 6 (a) we show the scaling of MITL losses with the AK gap distance, $d$, for three different lengths. The pulse has $I_{\rm peak}=20\,$MA/m and $\tau_{\rm peak}=100\,$ns. We observe that for small AK gap sizes, $Q_{\rm loss}/Q_{\rm tot}$ scales as a power law $\sim 1/d$ while for larger gap sizes that scaling does not hold. In practical applications, small gap sizes are usually preferred and then the $\sim 1/d$ scaling can be used.
@@
Fig. 6 (b) shows how MITL losses scale with peak current, $I_{\rm peak}$, for three (fixed) MITL lengths, pulse length $\tau_{\rm peak}=100\,$ns, and $d=5\,$mm. The fractional losses exhibit a maximum at a certain value of $I_{\rm peak}$, vanish for small currents, and decrease for large peak currents. We can understand this behavior in the following way. On the one hand, smaller peak currents have smaller losses due to the shorter time of uninsulated SCL emission: regardless of the drive current amplitude, SCL emission turns on only after the electric field exceeds some fixed threshold value. If the peak current is too low, that threshold value cannot be exceeded since for a fixed $\tau_{\rm peak}$, $E\sim I_{\rm peak}$. On the other hand, large peak currents induce larger diamagnetic effects, which lower the losses. Obviously, peak currents around the maximum of the curve should be avoided in MITL designs. A common scaling of MITL losses for large peak currents emerges in the form of $\sim I_{\rm peak}^{-0.8}$, which we leave as an empirical observation.
@@
In Fig. 6 (c) we show the scaling of MITL losses with MITL length for fixed $I_{\rm peak}=20\,$MA/m and $\tau_{\rm peak}=100\,$ns, for three different AK gap sizes. At shorter lengths, the current losses drop due to the decreased time of uninsulated SCL emission. In the limit of (infinitely) short MITLs, we converge to the losses predicted by the Child-Langmuir law in vacuum fields (not shown; see also the discussion of Fig. 5). Similarly, we can understand the increased current losses in longer MITLs as increased time of uninsulated SCL emission. Recall that magnetic insulation cannot happen in pure TEM fields, i.e., for which $E=cB$: in such fields, electrons perform a figure-8 motion [42], not a cyclotron motion. Our estimates have shown that the size of such a figure-8 orbit for typical fields of interest is much larger than the AK gap, i.e., such electrons are uninsulated. Magnetic insulation can occur after reflection from the load because the relation $E=cB$ no longer holds due to the interference (superposition) of the incoming and outgoing (reflected) waves. In longer MITLs, reflection takes longer to happen, hence, longer time of uninsulation and increased losses. For the range of lengths shown, we again uncover an empirical scaling law for large $L$, however, different for the different gap sizes.
@@
It is worth commenting on the losses in infinitely long MITLs. Although magnetic insulation never occurs in such MITLs, we reiterate that some of the current limiting mechanisms already discussed in sections II and III still apply: (i) the Child-Langmuir law limits the amount of current grossing the AK gap; and (ii) the non-local diamagnetic effects further lower the amount of loss current; in fact, the latter are maximally manifested in this case.
@@
Figure 6 (d) shows scaling of MITL losses with pulse duration. This scaling is the strongest of the four examples shown in the figure, showing exponential or super-exponential dependence (notice the semi-log scale of the plot). The losses for the 111\,1m-long MITL fall to zero past about $\tau_{\rm peak}\gtrsim 300\,$ns. These losses can also be understood in terms of uninsulation time. First, uninsulation time is longer in longer MITLs (as discussed above), hence, increasing losses for larger $L$. Second, the relation $E\sim dI/dt$ implies lower fields for longer pulse lengths (for a fixed AK gap), hence, shorter uninsulation times for longer pulse lengths. (Strictly speaking, this relation only applies in infinitely short MITLs or equivalently, in the limit of $c\rightarrow\infty$, see also Ref. [43]. However, it approximately applies for the temporal spatial MITL lengths considered here, while also giving an intuitive way of discussing the observed dependencies.) No common empirical scaling law emerges in this case.